# A Comprehensive Evaluation of Semantic Relation Knowledge of Pretrained Language Models and Humans

### v1

@@ 8.6 Model Size Analysis
Table 6: Performance difference between the largest and the smallest model for four model families, across relations. The largest
Metric Avg. Type BERT RoBERTa ALBERT OPT
Soundness ($\Delta\mathcal{S}$) Micro +.$+.04$ +.$+.09$ +.$+.11$ +.$+.07$
Macro +.$+.06$ +.$+.07$ +.$+.14$ +.$+.08$
Completeness ($\Delta\mathcal{C}$) Micro +.$+.05$ +.$+.04$ +.$+.09$ +.$+.06$
Macro +.$+.06$ +.$+.04$ +.$+.12$ +.$+.08$
[As/S]ymmetry ($\Delta\mathcal{A}\mid\Delta\mathcal{M}$) Micro +.$+.05$ +.$+.06$ +.$+.12$ +.$+.10$
Macro +.$+.07$ +.$+.07$ +.$+.16$ +.$+.12$
Prototypicality ($\Delta\mathcal{S}$) Micro +.$+.05$ +.$+.07$ +.$+.09$ +.$+.06$
Macro +.$+.06$ +.$+.07$ +.$+.14$ +.$+.08$
Distinguishability ($\Delta$AuDC) n.a. $+1.23$ $+1.42$ $+4.36$ $+4.59$
@@
According to the scaling law, models with large sizes should outperform their smaller counterparts, so one should see only positive differences when model size increases. We examine whether this holds for our tasks. Table 6 presents the differences in performance between the largest model and the smallest model in each model family, for all metrics we consider [^22] The term “[As/S]ymmetry” refers to symmetry scores for symmetric relations and asymmetry scores for asymmetric relations.. Except for AuDC, all metrics are reported as two different averages. Macro differences are averaged over relations, whereas micro differences are averaged over individual scores per target word.
As all differences in all metrics shown in Table 6 are positive, the largest model outperforms the smallest model in the same family in every case. If there is a performance boost, however, its size cannot be predicted from the model size increase alone. The same increase in size (330 million) generally affords RoBERTa a higher degree of improvement than it does BERT. The comparison between OPT and ALBERT is also illustrative in this respect. Despite ALBERT’s smaller increase in model size (from 12 to 235 million) when compared to that of OPT (from 125 million to 65 billion), ALBERT always experiences a larger performance boost than OPT. When the performance differences are broken down into individual relations (results not shown here, but in Appendix D), we find that for most model families, antonymy is the relation that shows the highest improvement among all relations, ranging from 0.14 to 0.34. In contrast, the highest score reached for any other relation is only 0.16.
So far, we have only looked at comparisons of largest and smallest model in a family. We now look at smaller increases from one model to the next-larger model within its model family. We count the number of cases when a smaller model performs better than its next-larger model for all metrics and all relations [^23] If a model family has $n$ members, we perform $n$-1 tests. A table of detailed results can be found in Appendix D.. If this happens, the function between model size and performance is non-monotonic.
OPT is the family with the most non-monotonic behaviour out of the model families we consider: smaller models outperform the larger at least once for almost every relation and metric pair (except for hyponymy soundness, where we found no significance). Non-monotonic size-performance behaviour can also be observed for ALBERT in the following cases: all metrics for hypernymy, hyponymy completeness and asymemtry, holonymy soundness and prototypicality, and antonymy symmetry.
@@
To sum up, large models generally outperform small models. However, the models’ performance does not always increase monotonically with model size. Therefore, for the learning of semantic relations, model size is not all the matters.
@@ 8.7 Pretraining Task Analysis
In factual probing tasks, MLMs has been shown to outperform CLMs [49, 8, 42], as we have discussed in Section 2. We now verify whether this phenomenon also holds for semantic relations.
@@
We first calculate differences between pairs of best-performing MLMs and best-performing CLMs across model families, per metric. The results are presented in Table 7.
@@
Table 7: Difference between the best models pretrained on different tasks (the best MLM minus the best CLM). All within-metric differences are statistically significant.
Relation Soundness $\Delta\mathcal{S}$ Completeness $\Delta\mathcal{C}$ [As/S]ymmetry $\Delta\mathcal{A}\mid\Delta\mathcal{M}$ Prototypicality $\Delta\mathcal{S}$
HYP $+$.15 $+$.06 $+$.07 $+$.09
HPO $+$.20 $+$.09 $+$.09 n.a.
HOL $+$.12 $+$.07 $+$.05 $+$.11
MER $+$.18 $+$.13 $+$.14 n.a.
ANT $+$.10 $+$.10 $+$.10 $+$.07
SYN $+$.17 $+$.14 $+$.24 $+$.19
@@
As there are only positive performance differences in Table 7, it is never the case that the best CLM outperforms the best MLM. The best-performing MLM is always either RoBERTa-large (R2) or BERT-large (B2), whereas the best CLM is always either OPT-13B (O6), OPT-30B (O7), or OPT-66B (O8); different conditions produce different pairs. For AuDC, we find that the best-performing pair is established by OPT-66B and RoBERTa-large, with a numerical difference of 1.02 in favour of the MLM.
We want to point out that in all cases, the losing CLM model is larger than the winning MLM model. The smallest size difference observed between the pair of best MLM and best CLM (O6 and R2) is 12 billion parameters.
We also observe that the performance difference between MLM and CLM varies according to relations. The largest difference can be observed for synonymy, followed by meronymy, hyponymy, hypernymy, antonymy and holonym, in this order. This is exactly the opposite observation that we saw for model size, where antonymy profited most from larger sizes.
@@
To gather further evidence, we run every MLM model against every CLM model, and count the number of pairs where the MLM is significantly better. Table 8 shows the results.
Table 8: Number of MLM-CLM pairs where the MLM significantly outperforms the CLM, out of a total of 64.
Relation Soundness Completeness. [As/S]ymmetry Prototypicality
HYP 48 46 40 48
HPO 57 60 60 n.a.
HOL 44 41 29 41
MER 61 57 56 n.a.
ANT 49 47 47 50
SYN 61 62 63 62
The performance of the MLM is significantly better for at least 41 out of 64 MLM-CLM pairs, except for holonymy under asymmetry, where it is 29 pairs. For AuDC, as there is no statistical test, we report MLM-CLM pairs that yield numerical differences: there are 40 such pairs. Overall, this confirms the general superiority of MLMs over CLMs, particularly if we consider that in these pairs there are many small MLMs outperforming larger CLMs. The gap between MLMs and CLMs cannot be bridged by increases in model sizes, despite the tendency of larger models to learn semantic relations better.
@@ 8.8 Word Frequency Analysis
For every model, we use the rank correlation metric Spearman’s $\rho$ to compare word-frequency metrics that we derive independently from COCA, against all metrics except AuDC. We exclude AuDC from this analysis, as AuDC is a metric that is not lexically determined. As soundness, completeness, and prototypicality are summary statistics calculated across target words, we first need to recompile individual scores per target word. These scores can be obtained using Equations (11), (13) and (22). We then correlate the scores of each target word in these three metrics with the target word’s frequency.
Soundness and completeness are relations that involve relata sets. We therefore also need to consider the frequency of relata, not only of target words [^24] Prototypicality scores do not use relatum sets as gold standard. Thus, we only report correlation with target words for prototypicality.. We calculate the correlation between the scores against the average and maximum frequency of relata in each gold relatum set. We choose the average and maximum because we need to consider the relatum set as a whole, as this is how the sets are used in the calculation of these metrics.
Asymmetry and symmetry require special treatment because they are defined on tuples involving two words (the target word and the relatum). Agents might be unfairly advantaged in recognizing symmetry or asymmetry by word frequency effects in two cases: 1) if both words in the tuple are common and 2) if one word in the tuple is far more common than the other. We use two metrics: average frequency, which can guard against the first case, and absolute frequency difference, which can guard against the second. We first recalculate the symmetry and asymmetry scores per tuple, using Equations (15) and (17), and then correlate them with the average frequency and the absolute frequency difference.
This allows us to determine correlation for certain metric and relation combinations: For soundness and completeness, there are coefficients for all six relations. For symmetry, there are coefficients for the two symmetric relations. For asymmetry, there are coefficients for the four asymmetric relations. For prototypicality, there are coefficients for hypernymy, holonymy, antonymy, and synonymy. Consequently, the total number of coefficients is $16\times(2\times 6+1\times 2+1\times 4+1\times 4)=352$. The results are that correlations for all models have medians below 0.30 across all metrics and relations considered [^25] Details can be found in Appendix E.. Hinkle et al [25] regard correlations below 0.30 as negligible. We only conclude that there is a certain influence of word frequencies on the performance of semantic relation tasks, as was to be expected, but we believe that it is unlikely to be the defining factor in semantic relation learning.
@@ 9 Limitations
In our word frequency analysis above, we found little correlation between scores and word frequency, but we only examined unigram frequency of individual words. Unigram frequency cannot use any information of syntagmatic or paradigmatic relationships between two words. Using frequency-based metrics that are able to exploit such information may lead to a different conclusion. Particularly, more complex frequency measures that capture syntagmatic or paradigmatic relationships would be desirable. However, they require a precise definition of such measures and carefully controlled experiments, which are aspects beyond the scope of the present study.
Our evaluation also suffers from the fact that we only prompt with individual words and disregard subwords. The vocabulary of PLMs is a mixture of words and subwords, where frequent words remain as they are and less frequent words are split into subwords. Recent models like Llama [60] minimize their vocabulary size, instead including more subwords in their vocabulary. This may disadvantage them in our evaluation (and we even excluded Llama based on low vocabularly overlap). As a result, we were unable to gain knowledge about other CLMs, some of which are more widely used than OPT.
@@ 10 Conclusion
Current PLMs are commonly used for a wide range of tasks, so it is important to gain a comprehensive understanding of their linguistic abilities. This study focuses on the semantic relation knowledge of MLMs and CLMs. In particular, it explicates the gap in semantic relation knowledge between current PLMs and humans. Our contributions are as follows.
@@
    We presented a prompt-based probing evaluation methodology that covers six aspects of semantic relation knowledge, namely soundness, completeness, symmetry, asymmetry, prototypicality, and distinguishability. Two of these metrics are novel, namely those for prototypicality and for distinguishability. We also employed established metrics in a new context.
@@
    Using these evaluation methods, we conducted a comprehensive evaluation of the above-mentioned aspects for six semantic relations, five of which were never empirically tested in probing experiments before.
@@
Our experiment afforded far more conditions and distinctions than previous studies, including a human ceiling and a comparison of CLMs with MLMs. In this way, we arrive at a richer characterization of PLMs’ capabilities with semantic relations than was possible before. Our main result is that PLMs fall short of achieving human-level performance on the extensive semantic relation tasks defined here.
@@
We experimentally studied the prototypicality effect as displayed by humans for the six relations. We were able to confirm prototype effects for hypernymy and antonymy that have been experimentally studied in the literature. We also found a prototype effect for holonymy and synonymy that was not known before. One of our most imporantant findings with respect to prototypicality was that hyponymy and meronymy showed little prototypicality.
@@
We also found that when it comes to learning semantic relations, a large model size does not guarantee better performance in our tasks. The type of PLMs also matters. MLMs consistently outperformed CLMs across all metrics for all semantic relations. It therefore seems likely that the bidirectional context utilized by MLMs is a crucial factor in learning semantic relations.
@@
Out of all relations, antonymy is the one where both humans and models performed best, a result which is stable across all metrics. We also observed an antonymy bias operating in all models: while they were able to distinguish antonymy from non-antonymy, they often misrecognize non-antonymy as antonymy. Humans performed well with antonymy in both directions, as has been anticipated in previous work [12, 27, 9]. Since the antonymy bias appears across models, it may be attributed to some distributional characteristics of the antonymy relation which make it fundamentally different from the other relations. For example, although antonymy is mainly a paradigmatic relation, it also has a strong syntagmatic aspect: antonymy pairs commonly co-occur in conjunction structures such as “ascent and descent” and “either day or night”. While the other relations may also display a mixture of paradigmatic and syntagmatic features, the effect is certainly not as strong as for antonymy. In future work, it might be fruitful to study how the dual distributional nature of antonymy affects models’ recognition of it. Such insights might lead to better learning methods for the other relations.
@@
Our study aims to contribute towards a future where PLMs can better understand language. The fact that PLMs struggle to understand several aspects of semantic relations contradicts the superiority of PLMs as observed in many NLP tasks. This superiority is commonly attributed to the assumption that PLMs are able to efficiently encode general semantic and linguistic knowledge. Our work, which is a thorough investigation of this ability, showed that this is evidently not so. Some other explanation for the good performance should be sought. Our methodology is able to substantiate such doubts; in general, it captures effects not seen before. We therefore consider it a prism through which to see the truth more clearly.

### v4

@@ 8.7 Model Size Analysis
Table 6: Performance difference between the large and small variant of three model families, across relations.
Metric Avg. Type BERT RoBERTa Llama
Soundness ($\Delta\mathcal{S}$)
Micro +.04+.04+ .04 +.09+.09+ .09 +.01+.01+ .01
Macro +.05+.05+ .05 +.07+.07+ .07 +.02+.02+ .02
Completeness ($\Delta\mathcal{C}$)
Micro +.05+.05+ .05 +.04+.04+ .04 +.01+.01+ .01
Macro +.06+.06+ .06 +.04+.04+ .04 +.01+.01+ .01
Symmetry ($\Delta\mathcal{M}$)
Micro +.13+.13+ .13 +.11+.11+ .11 +.10+.10+ .10
Macro +.13+.13+ .13 +.11+.11+ .11 +.09+.09+ .09
Prototypicality ($\Delta\mathcal{P}$)
Micro +.05+.05+ .05 +.07+.07+ .07 +.00+.00+ .00
Macro +.05+.05+ .05 +.06+.06+ .06 +.00+.00+ .00
Distinguishability ($\Delta$AuDC) n.a. +1.22+1.22+ 1.22 +1.40+1.40+ 1.40 +0.32+0.32+ 0.32
@@
According to the scaling law [32], large-size models should outperform their smaller counterparts, so one should see a positive difference when model size increases. We examine whether this holds for our tasks. Table 6 presents the differences in performance between the large and small models in each model family, for all metrics we consider. Except for AuDC, all metrics are reported as two different averages. Macro differences are averaged over relations, whereas micro differences are averaged over individual scores per target word.
As all differences in all metrics shown in Table 6 are positive, the large model outperforms the small model in the same family in almost every case (except for Llama on prototypicality, where the differences are zero). However, if there is a performance boost, its magnitude cannot be directly inferred from the model size increase alone. The same increase in size (330 million, 110M to 340M for BERT and 125M to 355M for RoBERTa) generally affords RoBERTa a higher degree of improvement in soundness, prototypicality, and distinguishability than it affords BERT. Despite the smaller increase in model size of BERT and RoBERTa, when compared to that of Llama (72 billion), the former two MLMs consistently exhibit a greater performance boost than Llama, where the improvements are overall marginal.
When the performance differences are broken down into individual relations, we find that the relations that benefit from the increase in model size vary across model families [^18] Results are shown in Appendix D).. Antonymy shows the highest improvement among all relations for BERT, ranging from 0.14 to 0.19. For RoBERTa, the highest improvement is usually observed in hypernymy or synonymy, ranging between 0.08 to 0.12. In contrast, increasing the size of Llama does not consistently yield better performance. For hypernymy and antonymy, the smaller Llama-8B generally significantly outperforms the larger Llama-70B across almost all metrics [^19] Except for antonymy symmetry, the boost in scores is only 0.04.. For holonymy soundness and prototypicality, as well as hyponymy completeness, there is no significant improvement.
@@
To sum up, while large models typically outperform small models, we have seen enough counterexamples to conclude that this regularity does not necessarily or even generally hold. Therefore, model size is not all that matters for the learning of semantic relations.
@@ 8.8 Pretraining Task Analysis
In factual probing tasks, MLMs have been shown to outperform CLMs [52, 10, 45] whereas currently CLMs are used more wildly in a serious of tasks [71, 64, 18], as we have discussed in Section 2. We now verify which pretraining task leads to a better quality of semantic relations knowledge.
@@
We first calculate differences between pairs of the best-performing CLM and MLM per metric. The best-performing MLM is always either RoBERTa-large (R2) or BERT-large (B2), whereas the best CLM could be either Llama-8B(L1) or Llama-70B(L2); different conditions produce different pairs. The results are presented in Table 7.
@@
Table 7: Difference between the best models pretrained on different tasks (the best CLM minus the best MLM). Statistically significant within-metric differences are reported as values; non-significant differences are parenthesized.
Relation
Soundness
$\Delta\mathcal{S}$
Completeness
$\Delta\mathcal{C}$
Symmetry
$\Delta\mathcal{M}$
Prototypicality
$\Delta\mathcal{P}$
HYP $+$.07 (.00) n.a. $+$.02
HPO --.10 --.07 n.a. n.a.
HOL $+$.06 $+$.03 n.a. --.02
MER --.12 --.10 n.a. n.a.
ANT $+$.08 $+$.07 $+$.05 (.03)
SYN $+$.13 $+$.06 $+$.05 (.03)
@@
As there are positive, negative, and non-significant performance differences in Table 7, it is never the case that the best CLM always outperforms the best MLM. For AuDC, we find that the best-performing pair is Llama-70B (L2) and RoBERTa-large (R2), with a numerical difference of 2.5 in favour of the CLM.
For prototypicality, there is either no significant difference at all, or if there is one, it is small. Combining this observation with the fact that even the best performance on prototypicality is below 0.44 (achieved by Llama-8B for antonymy), we can conclude that both types of models do not perform well under this metric.
For soundness and completeness of hyponymy and meronymy, the best CLM significantly falls behind the best MLM by at least 0.07. However, for soundness and completeness of hypernymy and holonymy, which are the reverse relations of hyponymy and meronymy, the best CLM shows either comparable or significantly better performance than the best MLM. In other words, CLMs are more sensitive to relation directionality. This sensitivity aligns with the reversal curse observed in the previous literature [5], where CLMs only solve factual prompting tasks in one direction but not in the reverse direction. We want to point out that in all cases, the CLM model is larger than the MLM model. Even the smallest size difference observed between the pair of best MLM and best CLM (RoBERTa-large and Llama-8B) is more than 6 billion parameters. Our results further show that this sensitivity is so strong that, when the direction is reversed, CLMs’ performance drops below that of MLMs, which are tens of times smaller in size.
@@
To gather further evidence, we compare the performance of every CLM against every MLM. In other words, for every MLM-CLM pair (8 pairs in total), we determine if the MLM significantly outperforms the CLM. We then count the number of pairs where the MLM is significantly better. Table 8 shows the results.
Table 8: Number of MLM-CLM pairs where the MLM significantly outperforms the CLM, out of eight pairs.
Relation Soundness Completeness.
Symmetry
Prototypicality
HYP 0 2 n.a 0
HPO 2 8 n.a n.a.
HOL 0 0 n.a 1
MER 6 7 n.a n.a.
ANT 0 0 0 0
SYN 0 1 2 1
The significant advantage of MLM can be observed particularly in hyponymy and meronymy, where at least 2 out of 8 MLM-CLM pairs belong to such cases202020 For AuDC, as there is no statistical test, we compare model pairs numerically. We do not find any MLM outperforming a CLM.. For hyponymy and meronymy completeness, almost all CLMs remain below MLMs.
Hence, this confirms that there is no general superiority of CLMs over MLMs, and that CLMs present weakness in learning hyponymy and meronymy, which is hard to even be bridged by increasing model sizes.
@@ 8.9 Word Frequency Analysis
For every model, we use the rank correlation metric Spearman’s $\rho$ to compare word-frequency metrics that we derive independently from COCA [16], against all metrics except AuDC. We exclude AuDC from this analysis, as AuDC is a metric that is not lexically determined. As soundness, completeness, and prototypicality are summary statistics calculated across target words, we first need to recompile individual scores per target word. These scores can be obtained using Equations (10), (12) and (19). We then calculate correlations between the scores of these three metrics with word frequencies.
Soundness and completeness are relations that involve relatum sets. We therefore also need to consider the frequency of relata, not only of target words [^21] Prototypicality scores do not use relatum sets as gold standard. Thus, we only report correlation with target words for prototypicality.. We calculate the correlation between the scores against the average and maximum frequency of relata in each gold relatum set. We choose the average and maximum because we need to consider the relatum set as a whole, as this is how the sets are used in the calculation of these metrics.
Symmetry requires special treatment because it is defined on tuples involving two words (the target word and the relatum). Agents might be unfairly advantaged in recognizing symmetry by word frequency effects in two cases: 1) if both words in the tuple are common and 2) if one word in the tuple is far more common than the other. We use two frequencies: average frequency, which can guard against the first case, and absolute frequency difference, which can guard against the second. We first recalculate the symmetry scores per tuple, using Equations (14), and then correlate them with the average frequency and the absolute frequency difference.
This allows us to determine correlation for certain metric and relation combinations: For soundness and completeness, there are coefficients for all six relations. For symmetry, there are coefficients for the two symmetric relations. For prototypicality, there are coefficients for hypernymy, holonymy, antonymy, and synonymy. Consequently, the total number of coefficients is 6×(2×6+1×2+1×4)=606$\times$(2$\times$ 6+1$\times$ 2+1$\times$ 4)=606 × ( 2 × 6 + 1 × 2 + 1 × 4 ) = 60. The results are that correlations for all models have medians below 0.30 across all metrics and relations considered [^22] Details can be found in Appendix E.. Hinkle et al [29] regard correlations below 0.30 as negligible.
As the metrics used for frequency are based on unigram, we conclude that the unigram frequency is unlikely to be the decisive factor in semantic relation learning.
@@ 9 Limitations
We chose to perform our evaluation at the word level, rather than the subword level. This decision was made in order to ensure the comparability between CLMs and MLMs. The number of masks that MLMs need to fill has to be determined in advance. However, CLMs can predict a sequence of subwords of any length. The pre-defined numbers of masks may provide extra hints to MLMs, which may give an advantage to them and hence may work against comparability.
In our word frequency analysis above, we found little correlation between scores and word frequency, but we only examined the unigram frequency of individual words. Unigram frequency cannot use any information of syntagmatic or paradigmatic relationships between two words. Using frequency-based metrics that can exploit such information may lead to a different conclusion. Particularly, more complex frequency measures that capture syntagmatic or paradigmatic relationships would be desirable. However, they require a precise definition of such measures and carefully controlled experiments, which are aspects beyond the scope of the present study.
@@ 10 Conclusion
Current PLMs are commonly used for a wide range of tasks, so it is important to gain a comprehensive understanding of their linguistic abilities. This study focuses on the semantic relation knowledge of PLMs. In particular, it explicates the gap in semantic relation knowledge between current PLMs and humans. Our contributions are as follows.
@@
    We presented a prompt-based probing evaluation methodology that covers five aspects of semantic relation knowledge, namely soundness, completeness, symmetry, prototypicality, and distinguishability. Two of these metrics are novel, namely those for prototypicality and distinguishability. We also employed established metrics in a new context.
@@
    Using these evaluation methods, we conducted a comprehensive evaluation of the above-mentioned aspects for six nominal semantic relations, five of which were underexplored in probing experiments before.
@@
Our experiment afforded far more conditions and distinctions than previous studies, including a human ceiling and a comparison of CLMs with MLMs. As a result, we arrive at a richer characterization of PLMs’ capabilities with semantic relations than was possible before. Our main result is that PLMs fall short of achieving human-level performance on the extensive semantic relation tasks defined here.
@@
We experimentally studied the prototypicality effect as displayed by humans for the six relations. We were able to confirm prototype effects for hypernymy and antonymy that have been experimentally studied in the literature. We also found a prototype effect for holonymy and synonymy that was not known before. One of our most important findings for prototypicality was that hyponymy and meronymy showed little prototypicality.
@@
We also found that a large model size does not always guarantee better performance in learning semantic relations. The type of PLMs also matters. MLMs consistently outperformed CLMs for hyponymy and meronymy. Therefore, the bidirectional context utilized by MLMs is a crucial factor in learning these two semantic relations.
@@
We observed an antonymy bias: both humans and models performed best for antonymy out of all relations; the results are stable across all metrics. MLMs often misrecognize non-antonymy relations as antonymy. On the other hand, CLMs only misrecognize synonymy as antonymy. Humans performed well with antonymy in both directions. Since the antonymy bias appears across all the models evaluated, it may not be attributed to the pretraining tasks or the model sizes. One possible reason might be the distributional characteristics of the antonymy relation in the corpus, which make antonymy fundamentally different from the other relations. Antonym pairs commonly co-occur in conjunctive structures such as “ascent and descent” and “day or night”. The other relations might not present such strong co-occurrence tendencies. This contrast could explain the antonymy bias we observed in the present study. Future work includes studying how the co-occurrences of antonyms affect their recognition by models. Such insights should also lead to better learning methods for the other relations.
@@
Our study aims to contribute towards a future where PLMs can better understand language. The fact that PLMs struggle to understand several aspects of semantic relations contradicts the superiority of PLMs as observed in many NLP tasks. This superiority is commonly attributed to the assumption that PLMs are able to efficiently encode general semantic and linguistic knowledge. Such a strong assumption requires solid evaluation methodology to be able to formulate them scientifically and substantiate it. Our methodology is able to accomplish this. Through the comprehensive evaluation, our work puts doubts on this assumption. Some other explanations for the good performance should be sought. In general, it captures effects not seen before. We, therefore, consider it a prism through which to see the truth more clearly.